# Parametrically-controlled microwave-photonic interface for the fluxonium

### v2

@@ Appendix B Experimental setup
Figure 7: Simplified schematic of the measurement setup for the controlled release experiment. The qubit Rabi drive is directly output by an integrated FPGA system (Quantum Machines, OPX+). The pump tone for the $|e0\rangle\leftrightarrow|g1\rangle$ transition is applied by mixing drives from a microwave generator (SignalCore, SC5511A) as input LO and IQ tones from the FPGA system. The pump signal is first amplified and then sent into the dilution refrigerator (Oxford, Triton 300) along with the qubit Rabi drive through the sample’s charge line. The output signal from the sample goes through a circulator and is amplified by a traveling wave parametric amplifier (TWPA) at the mixing chamber (MXC). Above the MXC, the output signal goes through two additional circulators. At room temperature, the signal gets amplified before demodulated down to 50 MHz. We use the same input LO generator for output demodulation, so that the output signal is phased locked to the output LO. The demodulated signal goes through three 41-58 MHz band-pass filters before a final stage of amplification. The signal is then further demodulated by analog-to-digital converters (ADC) in the FPGA system.
Figure 7 shows the simplified schematic of the measurement setup used for the controlled-release measurement demonstrated in Section V. For this experiment, it is important to ensure that the photon release and photon capture processes are phase-locked. To ensure phase-locking of the release process, we apply the Rabi pulse on the qubit directly using the I channel output from the FPGA system, and modulate the pump tone addressing the $|e0\rangle\leftrightarrow|g1\rangle$ transition using the same FPGA system. To ensure phase-locking of the output signal, we use the pump signal generator to provide the local oscillator for output demodulation. The pump tone is amplified before sent into the dilution refrigerator. The pump tone and the qubit Rabi drive couple to the SNAIL-fluxonium through the charge line weakly coupled to the SNAIL’s antenna.
On the output side, the signal leaving the resonator is first routed through a circulator before being amplified by a TWPA. The signal goes through two additional circulators thermalized to the MXC. At room temperature, the output signal is amplified and then demodulated down to 50 MHz. We use the same input LO generator for output demodulation, so that the output signal is phase-locked to the output LO. The demodulated signal goes through 41-58 MHz band-pass filters to filter away any pump tones that leak to the readout line. After filtering, the signal gets amplified before being demodulated by the FPGA.
@@ Appendix C System Hamiltonian
$\hat{H}_{\text{3WM+d}}/\hbar=c_{3}^{s}(\phi_{f,\mathrm{zpf}}(\hat{f}^{\dagger}+\hat{f})+\phi_{s,\mathrm{zpf}}(\hat{s}^{\dagger}+\hat{s})\\ +\phi_{r,\mathrm{zpf}}(\hat{r}^{\dagger}+\hat{r}))^{3}+\epsilon(t)e^{-i\omega_{d}t}(\hat{s}^{\dagger}+\hat{s}).$ (7)
Going into the displaced frame, we introduce the displaced amplitude,
$\tilde{\xi}(t)=\xi(t)e^{-i\omega_{d}t}=\frac{\epsilon(t)e^{-i\omega_{d}t}}{i\kappa_{s}/4+\Delta_{c}}\approx\frac{\epsilon(t)e^{-i\omega_{d}t}}{\Delta_{c}},$ (8)
@@
where $g(t)=6c_{3}^{s}\xi(t)\prod\limits_{i=f,s,r}\phi_{i,\mathrm{zpf}}$ is the coupling rate of the $|e0\rangle\leftrightarrow|g1\rangle$ transition, and $\delta=\omega_{d}-(\omega_{r}-\omega_{f})$ is a small detuning between the pump frequency and the $|e0\rangle\leftrightarrow|g1\rangle$ transition frequency. In simulation, we find that computing the bare SNAIL phase matrix element squared $6c_{3}^{s}\xi(t)|\langle g1|\hat{\phi}_{s}^{2}|e0\rangle|$ already gives useful guidance for choosing system parameters.
@@
$\delta\hat{\rho}=-i[\hat{H}_{\text{eff}},\hat{\rho}]+\sum_{k=g0,e0,g1}[L_{k}\rho L_{k}-\frac{1}{2}\{L_{k}L_{k},\hat{\rho}\}].$ (12)
Here $\hat{\rho}$ is the system density matrix operator. $\hat{H}_{\text{sys}}$ is defined in equation (1) and $\hat{H}_{d}$ is the driving Hamiltonian. The jump operators $L_{g0}=\sqrt{\Gamma_{\downarrow}}|e0\rangle\langle g0|,L_{e0}=\sqrt{\Gamma_{\uparrow}}|g0\rangle\langle e0|,$ and $L_{g1}=\sqrt{\kappa}|g0\rangle\langle g1|$ describe the main depolarization paths of the quantum system.
@@ Appendix E Shaping wave packets
In Fig. 9(a), we show both numerically and analytically computed coupling strength for producing a hyperbolic secant wave packet with $\gamma_{\text{ph}}/2\pi=0.2$ MHz. Both methods yield coupling strength $g(t)$ consistent with each other. In Fig. 9(b), we show the time evolution of the fluxonium population in $|e0\rangle$ under our simulated pump tone without state initialization. The fluxonium population decays from thermal equilibrium and saturates at $8\%$, consistent with our reset fidelity as well as numerical fit using the master equation. This indicates that our applied pump pulse amplitude evacuates a fluxonium excitation, the fidelity of which is limited by our device’s $T_{1}$ decay during state initialization and controlled release processes due to our low decay rate $\kappa$. By simply designing a higher $\kappa$, one can achieve more efficient controlled release process without additional tuning of design parameters.

### v3

@@ Appendix B Experimental setup
Figure 8: Simplified schematic of the measurement setup for (a) two-tone spectroscopy and (b) controlled release experiment. In (a), both the readout tone and the pump tone for driving the $|e0\rangle\leftrightarrow|g1\rangle$ transition are generated by mixing local oscillator (LO) from a microwave generator (SignalCore, SC5511A) and IQ tones from an integrated FPGA system (Quantum Machines, OPX+). The pump signal is first amplified and then sent into the dilution refrigerator (Oxford, Triton 500) through the sample’s charge line. The readout tone transmits through the readout line. The output signal from the sample goes through a circulator and is amplified by a traveling wave parametric amplifier (TWPA) at the mixing chamber (MXC). Above the MXC, the output signal goes through two additional circulators. At room temperature, the signal gets amplified before demodulated down to 50 MHz. The demodulated signal goes through three 41-58 MHz band-pass filters before a final stage of amplification. The signal is then further demodulated by analog-to-digital converters (ADC) in the FPGA system. In (b), the qubit Rabi drive is directly output by the FPGA system. The qubit Rabi drive and the pump tone are combined and sent through the charge drive line. No readout tone is applied to the system, as we are directly reading out the photon released from the readout resonator. We use the same pump LO generator for output demodulation, so that the output signal is phased locked to the output LO.
Figure 8 shows the simplified schematic of the measurement setup used for the two-tone spectroscopy measurement in Section II and the controlled-release experiment demonstrated in Section V. To perform two-tone spectroscopy, a readout tone is sent to the sample through the readout line while the amplified pump tone is sent to the qubit through the charge line weakly-coupled to the SNAIL’s antenna. On the output side, the transmitted readout tone is first routed through a circulator before being amplified by a TWPA. The signal goes through two additional circulators thermalized to the MXC. At room temperature, the output signal is amplified and then demodulated down to 50 MHz by the readout LO. The demodulated signal goes through 41-58 MHz band-pass filters to filter away any pump tones that leak to the readout line. After filtering, the signal gets amplified before being demodulated and digitalized by the FPGA for obtaining the I and Q components of the transmitted readout tone.
For the controlled release experiment, it is important to ensure that the photon release and photon capture processes are phase-locked. To ensure phase-locking of the release process, we apply the Rabi pulse on the qubit directly using the I channel output from the FPGA system, and modulate the pump tone addressing the $|e0\rangle\leftrightarrow|g1\rangle$ transition using the same FPGA system. To ensure phase-locking of the output signal, we use the pump signal generator to provide the LO for output demodulation. No readout tone is applied to the sample, since the photon we capture is emitted through conversion of the fluxonium excitation into a readout resonator photon.
@@ Appendix C System Hamiltonian
$\hat{H}_{\text{3WM+d}}/\hbar=c_{3}^{s}[\phi_{f,\mathrm{zpf}}(\hat{f}^{\dagger}+\hat{f})+\phi_{s,\mathrm{zpf}}(\hat{s}^{\dagger}+\hat{s})\\ +\phi_{r,\mathrm{zpf}}(\hat{r}^{\dagger}+\hat{r})]^{3}+2\epsilon(t)\cos{(\omega_{d}t)}(\hat{s}^{\dagger}+\hat{s}),$ (8)
where $\epsilon(t)$ is the time-dependent pump amplitude. Going into the displaced frame, we introduce the displaced amplitude,
$\tilde{\xi}(t)=\xi(t)e^{-i\omega_{d}t}=\frac{\epsilon(t)e^{-i\omega_{d}t}}{i\kappa_{s}/2+\Delta_{c}}\approx\frac{\epsilon(t)e^{-i\omega_{d}t}}{\Delta_{c}},$ (9)
@@
where $g(t)=6c_{3}^{s}\xi(t)\phi_{f,\mathrm{zpf}}\phi_{s,\mathrm{zpf}}\phi_{r,\mathrm{zpf}}$ is the coupling rate of the $|e0\rangle\leftrightarrow|g1\rangle$ transition, and $\delta=\omega_{d}-(\omega_{r}-\omega_{f})$ is a small detuning between the pump frequency and the $|e0\rangle\leftrightarrow|g1\rangle$ transition frequency. In simulation, we find that computing the bare SNAIL phase matrix element squared $6c_{3}^{s}\xi(t)|\langle g1|\hat{\phi}_{s}^{2}|e0\rangle|$ already gives useful guidance for choosing system parameters.
@@
In Fig.9, we show two-tone spectroscopy data of the $|e0\rangle\leftrightarrow|g1\rangle$ transition under 300 ns pump pulse as a function of pump power at the chip. In the weak coupling regime, the $|e0\rangle\leftrightarrow|g1\rangle$ transition frequency is centered at 6.16455 GHz. In the strong coupling regime, the $|e0\rangle\leftrightarrow|g1\rangle$ transition frequency is only shifted by 200 kHz due to the stronger pump tone and is centered at 6.16435 GHz. One can estimate the number of photons sent into the SNAIL via the formula [30]
$|\xi|^{2}=\bar{n}=2\kappa_{\mathrm{ext}}|a_{\mathrm{in}}|^{2}/(\kappa_{\mathrm{tot}}+4\Delta^{2}),$ (12)
where $|a_{\mathrm{in}}|^{2}$ is the incoming photon flux in unit of photon/s. Assuming internal and external quality factors of $Q_{\mathrm{in}}\approx 5\times 10^{4},Q_{\mathrm{ext}}\approx 5\times 10^{5}$ and an expected frequency of $\omega_{s}/2\pi=6.93$ GHz for the SNAIL, we estimate $|\xi|^{2}\approx 0.088$ in the weak coupling regime and $|\xi|^{2}\approx 0.86$ in the strong coupling regime, both of which are below single photon level.
@@
$\delta_{t}\hat{\rho}=-i[\hat{H}_{\text{eff}},\hat{\rho}]+\sum_{k=g0,e0,g1}[L_{k}\rho L_{k}-\frac{1}{2}\{L_{k}L_{k},\hat{\rho}\}].$ (14)
Here $\hat{\rho}$ is the system density operator. $\hat{H}_{\text{sys}}$ is defined in equation (1) and $\hat{H}_{d}$ is the driving Hamiltonian. The jump operators $L_{g0}=\sqrt{\Gamma_{\downarrow}}|e0\rangle\langle g0|,L_{e0}=\sqrt{\Gamma_{\uparrow}}|g0\rangle\langle e0|,$ and $L_{g1}=\sqrt{\kappa}|g0\rangle\langle g1|$ describe the main depolarization paths of the quantum system.
@@
Figure 9: Two-tone spectroscopy data of $|e0\rangle\leftrightarrow|g1\rangle$ transition under 300 ns pump pulse as a function of pump power estimated at the chip. Dashed lines label the different power levels we use to realize weak (red) and strong (blue) coupling demonstrated in Fig.2.
@@ Appendix E Shaping wave packets
In Fig. 11(a), we show both numerically and analytically computed coupling strength for producing a hyperbolic secant wave packet with $\gamma_{\text{ph}}/2\pi=0.2$ MHz. Both methods yield coupling strength $g(t)$ consistent with each other. In Fig. 11(b), we show the time evolution of the fluxonium population in $|e0\rangle$ under our simulated pump tone without state initialization. The fluxonium population decays from thermal equilibrium and saturates at $8\%$, consistent with our reset fidelity in the strong coupling regime as well as numerical fit using the master equation. This indicates that our applied pump pulse amplitude evacuates a fluxonium excitation, the fidelity of which is limited by our device’s $T_{1}$ decay during state initialization and controlled release processes due to our low decay rate $\kappa$. By simply designing a higher $\kappa$, one can achieve more efficient controlled release process without additional tuning of design parameters.
Figure 12: (a) Numerically simulated field amplitude squared as a function of pulse duration for the $(|0\rangle+|1\rangle)/\sqrt{2}$ state generated in Fig. 4(c). Solid line includes both state preparation error and conversion inefficiency. Dotted line corrects for state preparation error. Dashed line denotes perfect conversion efficiency. (b) Numerically simulated field amplitude squared as a function of pulse duration similar to that in Fig. 12(a) but with a higher resonator decay rate $\kappa/2\pi=4$ MHz, which allows for faster qubit-photon conversion.
[…]
where $\hat{r}_{0}(t)$ denotes outgoing field amplitude under perfect conversion. The total efficiency is simulated to be $\sim 79\%$ including state preparation error. If correcting for state preparation error, the conversion efficiency rises to $\sim 86\%$, the loss of which is mainly due to qubit depolarization during the controlled release process. However, such source of error can be easily suppressed by designing a larger resonator decay rate, which allows for a faster conversion process. With a resonator decay rate of $\kappa/2\pi=4\mathrm{MHz}$, one can reach a conversion efficiency of $\eta\approx 99\%$ under perfect state initialization (Fig. 12(b)). If we include state preparation error assuming $\sim 99.8\%$ initialization efficiency (Fig. 10(d)), the conversion efficiency is only compromised by an extra loss of $\sim 0.4\%$.
Finally, a flying qubit shaped into a time-symmetric profile can be absorbed with unit probability by time-reversing the release process on a second qubit $g_{2}(t)=g_{1}(-t)$ [2]. We quantify the symmetry of our generated wave packet by computing the following symmetry parameter [17]
[…]
which yields $s=0.975$. By increasing the resonator decay rate, the symmetry parameter of the shaped wave packet can be further optimized by avoiding asymmetry due to the resonator rise time and qubit depolarization.